OpenStudy (anonymous):

Which ordered pair is a solution of the equation 2x − y = 9 (-4,1) (-2,5) (5,1) (6,-3)

OpenStudy (anonymous):

(6,-3)

OpenStudy (anonymous):

2(6)-3=9

OpenStudy (conqueror):

C

OpenStudy (conqueror):

(5, 1)

OpenStudy (anonymous):

@SkylarLee thats not right Correct answer is (5,1)

OpenStudy (conqueror):

@mathmale Am I correct?

OpenStudy (anonymous):

I used it in plato and it did come out correct. me and @tk23 are doing the same thing

OpenStudy (conqueror):

Let's see what mathmale says

OpenStudy (anonymous):

its some glitch and plato with that unit it did that

OpenStudy (anonymous):

(5,1) https://www.desmos.com/calculator

OpenStudy (anonymous):

httsp/:platoadvance.gle.com

OpenStudy (anonymous):

^ it said it was a glitch ._.

OpenStudy (anonymous):

its not correct but it uses four different answersany answer he picked woulld have been right in that unit bc plato is glitching do you understand lol @Kemist

OpenStudy (anonymous):

(5,1) is correct.

OpenStudy (conqueror):

Look at that graph and tell me what you see

OpenStudy (mathmale):

Instead of asking me (or someone else), "Am i correct?" why not ask, "I want to be able to check my own answers. Could you give me some pointers on how to do that?"

OpenStudy (anonymous):

Well as long as he gets it right then I guess the answers doesn't matter :D

OpenStudy (conqueror):

@mathmale I just proved I'm right with that graph :P

OpenStudy (anonymous):

@Conqueror (5,1) is correct but if he would have put it in on plato in that unit it will mark it wrong even tho its the right answer for some reason plato is doing that i had to type in the wrong answers we dunno whats wrong >_<

OpenStudy (conqueror):

Don't use "plato"

OpenStudy (anonymous):

https://www.desmos.com/calculator Is great for these problems.

OpenStudy (mathmale):

What does the word "solution" mean when we're discussing "solution of the equation 2x − y = 9"?

OpenStudy (conqueror):

Indeed @Kemist

OpenStudy (mathmale):

Let's please cease this "I'm right!" stuff and concentrate on learning how to check your own work and how to help others check their own work.

OpenStudy (mathmale):

Who's willing and able to complete the sentence, "a given ordered pair is a solution to the equation 2x − y = 9 if ...... " ?